OpenStudy (pink.1d):

Prove: cos θ - cos θ sin2 θ = cos3 θ. You must show all work

OpenStudy (pink.1d):

FAN AND MEDAL :)

OpenStudy (pink.1d):

@Preetha @phi @dan815 @hero @Kainui

OpenStudy (pink.1d):

@triciaal

OpenStudy (anonymous):

\[\cos \theta -\cos \theta \sin ^2\theta=\cos \theta \left( 1-\sin ^2\theta \right)=\cos \theta \cos ^2\theta=?\]

OpenStudy (anonymous):

i am not here to give you straight forward answer.

OpenStudy (pink.1d):

IDK HELP ME GIVE ME A HINT

OpenStudy (anonymous):

what is x*x^2=

OpenStudy (pink.1d):

x^3

OpenStudy (anonymous):

correct now you come to my hint

OpenStudy (pink.1d):

oh wow wait its

OpenStudy (thatonegirl_):

oh good luck, i studied this stuff for a week and got a 63 on my test xD

OpenStudy (pink.1d):

cosθ^3

OpenStudy (pink.1d):

@Thatonegirl_ 63? its out of 50

OpenStudy (anonymous):

it is \[\cos ^3\theta \]

OpenStudy (pink.1d):

thats what i said

OpenStudy (pink.1d):

@surjithayer

OpenStudy (anonymous):

it was only correcting you.

OpenStudy (pink.1d):

@surjithayer ok sorry for being mean i gave u a medal to make up for that

OpenStudy (thatonegirl_):

No I got a 63% lol